# An embedding-based distance for temporal graphs

### v1

@@ Introduction
We now proceed to evaluate the distance metrics we introduced, focusing in particular on their ability to capture the differences in topological and temporal properties of empirical temporal graphs. We carry out two evaluation tests using both empirical and synthetic temporal graph data. We use, in particular, empirical temporal graphs published by the SocioPatterns collaboration (sociopatterns.org), that describe time-resolved, close-range proximity interactions of humans and animals in a variety of real-world environments. These data are known to exhibit rich multi-scale network structures, complex temporal activity patterns, correlations between these two, and other complex network features. Hence, they provide a natural benchmark to assess the sensitivity of our distance definitions to specific properties of real-world temporal graphs.
@@
  • In the first evaluation test, we consider structurally different synthetic temporal graphs of different sizes, and we cluster them using our temporal graph distance $d_{\rm u}$, checking how well the clusters we find match the graph generating methods we used, irrespective of graph size. We then repeat the same task on empirical labeled data of human proximity in school settings.
@@
  • In the second evaluation test, we use the empirical data to elucidate which of the many properties of real temporal graphs are discriminated by the distances we introduced. We proceed as follows: given an empirical temporal graph, we apply a set of randomization operations, where each operation is designed to preserve specific properties of the original data (e.g., the activity time series of each node). We then attempt to discriminate the randomized temporal graphs from one another using our distance definitions and quantify the performance of such discrimination task. We test both $d_{\rm m}$ and $d_{\rm u}$; however this evaluation strategy, by design, compares temporal graphs with the same number of nodes.
Discriminating between temporal graph classes
We define four families of temporal graphs as follows: first, we generate static random graphs of different sizes using four standard generative models; then, we turn these graphs into temporal graphs by assigning to each edge of the static graph the activity time series (i.e. the set of timestamps in which that edge was active) of a randomly chosen edge of the Conference empirical data set (the largest among the empirical data sets we use). The four generative models we use are: i) the Erdős-Renyi (ER) model [57]; ii) the stochastic block model (SBM), capable of generating graphs with a known community structure [58]; iii) the configuration model (CM), yielding graphs with an arbitrary degree distribution [59]; iv) the geometric model (GM), which generates edges based on proximity in a latent space [60]. Unlike the first three, the latter model yields graphs with a high clustering coefficient. Details on these generative models are provided in the Methods section.
[…]
@@
For each model, we generate $250$ graphs with size uniformly distributed between $n=200$ and $n=1800$ and constant average degree. We then deploy our distance definition $d_{\rm u}$ to compute a pair-wise distance matrix between all these temporal graphs. We use this matrix as an input for an unsupervised clustering task and check whether this clustering can discriminate the four groups of temporal graphs generated. The unsupervised clustering algorithm only takes the distance matrix and the number of clusters $k$ (here $k=4$) as inputs. Here, we first perform a spectral embedding using non-negative matrix factorization (NMF) [61] extracting $k$ non-negative components and then applying the k-means clustering algorithm with $k$ classes [62]. We note that successively embedding and clustering is a standard approach, and using NMF is most suited here since the distance matrix is non-negative. Finally, we compare the original temporal graph labels with the partition labels yielded by this unsupervised clustering informed by our temporal graph distances: we quantify the match between the true partition and the inferred one using the normalized mutual information (NMI), which yields a performance metric ranging between 00 (random guess) and $1$ (perfect reconstruction).
@@
Figure 2A shows the normalized mutual information (NMI) between the known temporal graph classes (associated with the four generative models) and the discovered cluster labels as a function of the embedding dimension $d$. The discrimination performance is low for small values of $d$, but for $d\gtrsim 8$ the accuracy is high and becomes insensitive to the specific $d$ value. The inset in Figure 2A shows the output of UMAP algorithm [63] in two dimensions when applied to the vectors $\bm{\lambda}\in\mathbb{R}^{32}$ used in the distance $d_{\rm u}$. Each symbol corresponds to one temporal graph, the color and marker shape indicate its generative model, and the marker size is proportional to the number of graph nodes. We observe that graph size does not appear to affect distances systematically, and that the $\bm{\lambda}$ vectors used by the distance $d_{\rm u}$ allow us to discriminate all four generative models. Indeed, we have verified that the distances $d_{\rm u}$ between the graphs and the distances between their respective UMAP embedded vectors are strongly related, with a highly significant Spearman correlation coefficient of $\sim 0.92$.
[…]
@@
Here, we consider $9$ empirical SocioPatterns data sets, whose basic properties are summarized in Table 1 of the Methods section. Following Ref. [64], we select $6$ types of randomization operations $R_{i}$ ($i=1,\ldots,6$), described in detail in the Methods section. In all cases, we preserve the number of nodes and of snapshots of the original temporal graph.
@@
Figure 3: Distance-based graph clustering for ensembles of graphs generated according to different randomization techniques. Left panel: results for the matched distance $d_{\rm m}$ of (3). Right panel: results for the unmatched distance $d_{\rm u}$ of (4). Within each panel, each matrix corresponds to one of the $9$ SocioPatterns temporal graphs described in Table 1. The rows and columns of each matrix correspond to the same set of $6$ randomization techniques we used, described in the Methods sections, and based on the properties they preserve. Random: preserves the number of temporal edges; Random delta: preserves the number of temporal edges and contact duration distribution; Active snapshot: preserves the number of edges at each time-step and activity pattern of each node; Time: preserves the aggregated graph structure; Sequence; preserves the structure of each snapshot; Weighted degree: preserves the nodes’ cumulative interaction time. For each pair of randomizations, we infer the randomization method of each temporal graph via an unsupervised distance-based clustering algorithm, and we compare the inferred randomization method with the known true one. Each matrix entry reports (value and color coding) the accuracy of the inferred labels (randomization methods), quantified as the normalized mutual information (NMI) between the inferred and true labels.
@@
To quantify the discriminative performance of our temporal graph distance, we proceed as follows, separately for each empirical temporal graph: given the empirical temporal graph and a pair of randomization operations ($R_{i}$, $R_{j}$), we generate a set of $250$ realizations using $R_{i}$, labeled as 00, and a set of $250$ realizations using $R_{j}$, labeled as $1$, for a total of $500$ temporal graphs. We then compute the matrix of distances between these $500$ graphs and use it to cluster them in $2$ clusters, following the same procedure used for Fig. 2a. Figure 3 reports the NMI values between the real labels and the unsupervised algorithm result for each pair of randomizations ($R_{i}$, $R_{j}$) and both the matched distance $d_{\rm m}$ (left panel) and the unmatched distance $d_{\rm u}$ (right panel). Within each panel, each matrix refers to one empirical temporal graph, and its entries are the NMI values described above. A high NMI entry indicates that our distance discriminates well the temporal graphs generated according to the two randomization procedures in the corresponding row and column of the matrix entry, i.e., that the distance is sensitive to the properties preserved by either $R_{i}$ or $R_{j}$. Low NMI values may either correspond to an inability of the distance to capture those properties or simply reflect that the empirical temporal graph is statistically random as far as those properties are concerned, as observed in particular for several pairs of randomizations in the Baboons dataset. Thus, it is enough to find one empirical data set for which the distance can distinguish between a pair of randomizations to conclude that the distance is indeed sensitive to the corresponding graph properties.
@@ Conclusions
We have here introduced a novel definition of distance between pairs of temporal graphs. This definition entails two steps: first, the embedding of the temporal graphs into an Euclidean space, and then the definition of a distance measure in the embedded space. In the first step, we have considered an embedding that exploits time-respecting random walks on the temporal graphs. Indeed, such walks are known to depend on, and encode important structural and temporal properties of these graphs [65, 9, 55, 40]. In the second step, we have proposed two possibilities for the distance definition. On the one hand, if a mapping is known between the nodes of the graphs to be compared, we consider a distance definition that leverages such mapping. On the other hand, we put forward in the more general case a definition that makes it possible to compare graphs with arbitrarily different sizes (numbers of nodes). Note that in both cases, as the size of the embedding matrix does not depend on the number of temporal graph snapshots, it is possible to embed temporal graphs with different lengths in the same embedding space and thus to compute a distance between them. The evaluation of our pipeline on both synthetic and empirical data sets has shown that the proposed distance can capture structural differences (e.g., degree distribution, clustering coefficient, or presence of communities) as well as temporal differences (e.g., burstiness or node activity patterns) in the temporal graphs being compared.
@@
The methodology we have developed is actually very general and customizable, as several other choices could be made in both steps of the method, namely the choices of the embedding method and of the distance. Our choices are based on two main motivations: i) a theoretical one, given the importance of time-respecting random walks in encoding information on temporal graphs, and thus in generating distributed representations of temporal graphs [65, 9, 55]; ii) a numerical one, which led us to choose the EDRep algorithm of Ref. [56] because of its efficiency and scalability and the possibility to obtain definitions of distance that can as well be efficiently computed. Moreover, the results obtained from our evaluation procedures support these choices a posteriori.
@@
Our work entails some limitations worth mentioning. First, our evaluation strategy has focused on a limited set of temporal graphs, namely on graphs obtained from data describing proximity events between individuals. These graphs are however known to feature a broad variety of representative complex structural and temporal patterns such as temporally changing community structures, burstiness, broad distributions of event durations, etc [12, 14]. Second, we have considered only the cases either of fully matched graphs (bijection between the sets of nodes) or of no known matching. An interesting intermediate situation would be the one of partially matched graphs, in which only a subset of nodes is matched between the two graphs. Tackling this challenging setting would form an interesting extension of this work.
@@
Defining a distance between temporal graphs opens the door to a wealth of potential applications, of which our evaluation procedures are only possible examples. Such a distance can, on the one hand, support data analysis of temporal graph data by extending the procedure of Fig. 3 to a large set of randomizations [47]. Indeed, as discussed, the inability to distinguish pairs of randomizations of a given temporal graph can be regarded as a statistical property of the graph itself. On the other hand, computing a distance between graphs can provide a crucial model validation tool: many generative models for temporal graphs are proposed based on various mechanisms of interactions, and model validation is often simply based on the comparison of a set of statistical properties [38], while our distance definition makes comparison possible at a global level. In addition, the distance could be used in the process itself of generating synthetic temporal graphs, by informing deep generative models such as GANS [66], which have proved extremely powerful in generating realistic data [67, 68]. Generating synthetic contact data with realistic structural and temporal properties could then be used to simulate processes of interest in public health research [69, 70, 71, 72, 73, 74]. Finally, empirical (temporal) graphs are hard to anonymize [75, 76, 73] and, to minimize the risk of node re-identification, it might be needed to perform some perturbation operations on the graph before making it publicly available. However, such operations risk destroying some important patterns and information contained in the data. A way to tackle this issue could be to find a trade-off between re-identification risk and information loss, and the distance we have defined could be used to quantify the latter.

### v2

@@ Introduction
We now evaluate the distance definitions we introduced, particularly with respect to their sensitivity to important topological and temporal properties of empirical temporal graphs. We carry out three kinds of evaluation tests using both synthetic and empirical temporal graph data. We use, in particular, empirical data shared by the SocioPatterns collaboration (sociopatterns.org) describing time-resolved, close-range proximity interactions of humans and animals in a variety of real-world environments (see Table 1 of the Methods section). These data are known to exhibit rich multi-scale network structures, complex temporal activity patterns, correlations between these two, and other complex network features [61]. Hence, they provide a natural benchmark to assess the sensitivity of our distance definitions to several specific properties of real-world temporal graphs. It is important to note that these data sets are by no means an exhaustive sample of temporal networks. However, our goal here is to show that the distances we introduce are sensitive to complex properties of temporal networks that we know are present in these particular data sets. Therefore, the evaluations we outline below do not depend on testing our distances on a comprehensive set of temporal networks but rather rely on making tests using data that are well understood, have been and are widely used by the scientific community and are known to exhibit many important properties of temporal graph data.
  1. 1.
@@
    Discriminating between classes of temporal graphs. We consider topologically different synthetic temporal graphs of different sizes and we cluster them using our temporal graph distance $d_{\rm u}$, checking how well the clusters we find match the known classes of the generative method we used, irrespective of graph size. We then carry out the same test using empirical graphs on human proximity in school settings.
  2. 2.
    Detecting partial node relabeling. We consider a synthetic temporal graph, select a fraction of its nodes, shuffle their indices and compute the distance $d_{\rm m}$ between the original and the re-labeled graphs. We investigate the behavior of the distance $d_{\rm m}$ as a function of the fraction of re-labeled nodes.
  3. 3.
@@
    Discriminating between temporal graph randomizations. We use the empirical data to elucidate which of the many properties of real temporal graphs are discriminated by the distances we introduced. We proceed as follows: given an empirical temporal graph, we apply to it a set of randomization operations, where each operation is designed to preserve specific properties of the original data (e.g., the activity time series of each node). We then attempt to discriminate the randomized temporal graphs from one another using our distance definitions and quantify the performance of such discrimination tasks. We test both $d_{\rm m}$ and $d_{\rm u}$; however, by design, this evaluation strategy compares temporal graphs with the same number of nodes.
Discriminating between classes of temporal graphs
@@
To test the ability of our model to distinguish topological network features, we proceed as follows. We consider four random generative models for static graphs and generate $250$ instances from each model with size uniformly distributed between $n=200$ and $n=1800$ and constant average degree. The models we use are: i) the Erdős-Renyi (ER) model [62]; ii) the stochastic block model (SBM), capable of generating graphs with a known community structure [63]; iii) the configuration model (CM), yielding graphs with an arbitrary degree distribution [64]; iv) the geometric model (GM), which generates edges based on proximity in a latent space [65]. Unlike the first three, the latter model yields graphs with a high clustering coefficient. Details on these generative models are provided in the Methods section. We then turn these graphs into temporal graphs by assigning an activity time series to each edge of the static graph, i.e. a set of timestamps at which that edge is active. To do this, we copy the activity time series of a randomly chosen edge of an empirical temporal graph, namely the Conference data set, which has the largest number of nodes among the empirical data sets we use. Details on the properties of this data set are given in Table 1. We then deploy our distance definition $d_{\rm u}$ to compute a pairwise distance matrix between all these temporal graphs and use this matrix as an input for an unsupervised clustering algorithm to label each graph. The unsupervised clustering algorithm only takes the distance matrix and the number of clusters $k$ (here $k=4$) as inputs. Here, we first perform a spectral embedding of the distance matrix between graph pairs using non-negative matrix factorization (NMF) [66] extracting $k$ non-negative components and then applying the k-means clustering algorithm with $k$ classes [67]. We note that successively embedding and clustering is a standard approach, and using NMF is most suited here since the distance matrix is non-negative. Finally, we compare the original temporal graph labels (i.e., the classes of our generative model) with the partition labels yielded by unsupervised clustering based on our distances: we quantify the match between the true partition and the inferred one using the normalized mutual information (NMI), which yields a performance metric ranging between 00 (random guess) and $1$ (perfect reconstruction).
Figure 2: Validation of the distances on graphs of varying size. Panel A: Accuracy of the distance-based clustering against the ground truth classes in terms of normalized mutual information (NMI) as a function of the embedding dimensionality $d$ used in the embedding step shown in Figure 1. The clustering task consists in recognizing four synthetic temporal graph classes with an unsupervised algorithm based on the distance $d_{\rm u}$. The synthetic graphs classes are obtained by generating a graph from either of four models (stochastic block model (SBM), configuration model (CM), Erdős-Renyi (ER), and geometric model (GM)) with constant degree equal to $4.8$, and the temporal component is obtained by sampling the edge activity of an empirical graph, as detailed in the main text. Inset of panel A: Scatter plot of UMAP dimensionality reduction in two dimensions of the vectors $\bm{\lambda}$ appearing in the definition of $d_{\rm u}$ given in (4), with here $d=32$. Each point refers to a temporal graph; the color and marker code refer to the generative model of its static component, while the marker size is proportional to the number of nodes. Panel B: Each point is the $2$-dimensional UMAP embedding of the vector $\bm{\lambda}$ of the temporal graph obtained selecting two classes and a day of interaction for all possible $(c_{2},c_{2},day)$ triplets in the SocioPatterns data sets describing proximity temporal graphs in schools. The color and marker code are assigned according to the class labels: the primary school classes form a group on their own and the other three groups (Other-Other, Bio-Other, Bio-Bio) belong to the high school data sets, where Bio are the biology classes, and Other the remaining ones. In all cases, we use a temporal resolution equal to $10$ minutes to create the temporal graphs from the SocioPatterns data. Panel C: normalized matched distance between a temporal graph and itself upon partial node re-labeling, as a function of the fraction $\alpha$ of re-labeled nodes. Each plot considers one of the generative models used in panel A, with the same nomenclature used for the title. The – barely visible – shadow line is the standard deviation of the distance value across $25$ randomizations of the partial re-labeling.
@@
Figure 2A shows the normalized mutual information (NMI) between the known temporal graph classes (associated with the four generative models) and the discovered cluster labels, as a function of the embedding dimension $d$. The discrimination performance is low for small values of $d$, but for $d\gtrsim 8$ the accuracy is high and becomes insensitive to the specific value of $d$. The inset in Figure 2A shows the output of UMAP algorithm [68] in two dimensions when applied to the vectors $\bm{\lambda}\in\mathbb{R}^{32}$ used in the distance $d_{\rm u}$. Each symbol corresponds to one temporal graph, the color and marker shape indicate its generative model and the marker size is proportional to the number of graph nodes. We observe that graph size does not appear to affect distances systematically and that the $\bm{\lambda}$ vectors used by the distance $d_{\rm u}$ allow us to discriminate all four generative models. Indeed, we have verified that the distances $d_{\rm u}$ between the graphs and the distances between their respective UMAP embedded vectors are strongly correlated, with a highly significant Spearman correlation coefficient of $\sim 0.92$.
We now carry out a test using the SocioPatterns empirical data sets describing time-resolved, close-range proximity interactions in schools over multiple days. These data sets are Primary school, High school 1, High school 2, High school 3. In these graphs, the nodes can be grouped according to school classes. In the primary school data set, there are ten classes, subdivided into two sections and five grades. In the high school data sets, the classes are defined by the main topics studied by the students: these topics can be maths and physics, physics and chemistry, or biology. The $2011$ data set is composed of three classes on physics and chemistry; the $2012$ data set has two additional classes on physics and maths; the $2013$ data set contains another physics and maths class and three biology classes. The metadata grouping nodes into classes are provided alongside the temporal networks data sets. We then assign one label to each class: the label is Primary for all primary school classes, while for the high school data sets, we only make the distinction between biology classes, labeled as Bio, and all others, labeled as Other.
Given a pair of classes $c_{1}$, $c_{2}$ from the same data set, we construct triplets $(c_{1},c_{2},day)$ inducing sub-temporal graphs formed by all the interactions on that day between the students of the two classes. For each sub-temporal graph, we compute the vector $\bm{\lambda}$ of Equation (4) and further decrease its dimensionality to $d=2$ using UMAP. We find that the Spearman correlation between the distances $d_{\rm u}$ and the distances between the UMAP embedding vectors is very high, namely equals $0.88$. Finally, we plot the resulting embedding in Fig. 2B, in which each point corresponds to a triplet. The color and marker coding denote the types of the two interacting classes. The plot shows that the distance can disentangle the networks formed from primary school data from those containing high school data. Moreover, a less neat but still visible distinction exists in the embeddings of (Bio-Other) with respect to the other triplets, likely because biology students spent more time in the labs and had different interaction patterns. This result confirms the ability of our distance to capture information on the temporal graph at multiple levels.
Detecting partial node relabeling
One required property for matched distance is to distinguish between graphs that differ only upon a node re-labeling. We work with the synthetic temporal graphs defined in section Discriminating between classes of temporal graphs to evaluate our distance in accomplishing this task. For each random graph model, we i) generate an instance of the random graph with $n=1000$ nodes; ii) randomly select a fraction $\alpha$ of nodes and shuffle their labels; iii) compute the distance between the original graph and the partially re-labeled one. This is repeated for $25$ different realizations of the partial re-labeling. Figure 2C shows the distance $d_{\rm m}$ (rescaled by the graph size) as a function of $\alpha$. The plots confirm for all the considered generative models that the distance depends on partial node re-labeling, with a positive distance as soon as $\alpha>0$, and increases with the fraction $\alpha$ of relabeled nodes. Notably, the plot suggests a high discriminative power even for small $\alpha$ values, as the distance increases rapidly with $\alpha$.
@@
Let us now consider $9$ empirical SocioPatterns data sets, whose basic properties are summarized in Table 1 of the Methods section. Following Ref. [69], we select $6$ types of randomization operations $R_{i}$ ($i=1,\ldots,6$), described in detail in the Methods section. In all cases, we preserve the number of nodes and the number of snapshots of the original temporal graph. These six randomizations are chosen because they allow us to inspect relevant temporal graph properties, including the interaction duration distribution, the node activation time series, the strength distribution in the aggregated graph, or the presence of a community structure. We can thus investigate graph properties both at the topological and temporal levels.
@@
Figure 3: Distance-based graph clustering for ensembles of temporal graphs generated according to different randomization techniques. Left panel: results for the matched distance $d_{\rm m}$ of (3). Right panel: results for the unmatched distance $d_{\rm u}$ of (4). Within each panel, each matrix corresponds to one of the $9$ SocioPatterns temporal graphs described in Table 1. The rows and columns of each matrix correspond to the same set of $6$ randomization techniques we used, described in the Methods sections. Each input graph is represented as a sequence of temporal edges $(i,j,t)$ as per Definition 1. The randomizations act on the temporal edges as follows. Random: preserves the number of temporal edges and randomizes the node and time indices; Random delta: preserves the number of temporal edges and interaction duration distribution; Active snapshot: preserves the number of edges at each time-step and the times at which each node is active, i.e., at which it has at least one neighbor. Time: preserves the aggregated weighted graph structure, i.e., the number of times each edge is active in the temporal graph; Sequence; preserves each snapshot’s adjacency matrix and randomizes the order in which they appear; Weighted degree: preserves the total number of temporal edges involving each node. For each pair of randomizations, we infer the randomization method of each temporal graph via an unsupervised distance-based clustering algorithm, and we compare the inferred randomization method with the known true one. Each matrix entry reports (value and color coding) the accuracy of the inferred labels (randomization methods), quantified as the normalized mutual information (NMI) between the inferred and true labels.
@@
To quantify the discriminative performance of our temporal graph distance, we proceed as follows, separately for each empirical temporal graph: given the empirical temporal graph and a pair of randomization operations ($R_{i}$, $R_{j}$), we generate a set of $250$ realizations using $R_{i}$, labeled as 00, and a set of $250$ realizations using $R_{j}$, labeled as $1$, for a total of $500$ temporal graphs. We then compute the matrix of distances between these $500$ graphs and use it to cluster them in $2$ clusters, following the same procedure used for Fig. 2A. Figure 3 reports the NMI values between the real labels and the unsupervised algorithm result for each pair of randomizations ($R_{i}$, $R_{j}$), and for both the matched distance $d_{\rm m}$ (left panel) and the unmatched distance $d_{\rm u}$ (right panel). Within each panel, each matrix refers to one empirical temporal graph, and its entries are the NMI values described above. A high NMI entry indicates that our distance discriminates well the temporal graphs generated according to the two randomization procedures in the corresponding row and column of the matrix entry, i.e., that the distance is sensitive to the properties preserved by either $R_{i}$ or $R_{j}$. Low NMI values may either correspond to an inability of the distance to capture those properties or simply reflect that the empirical temporal graph is statistically random as far as those properties are concerned, as observed in particular for several pairs of randomizations in the Baboons data set. Thus, it is enough to find one empirical data set for which the distance can distinguish between a pair of randomizations to conclude that the distance is indeed sensitive to the corresponding graph properties.
@@ Conclusions
We introduced a novel definition of distance between pairs of temporal graphs. This definition entails two steps. First, the temporal graphs are embedded in Euclidean space, and then a distance is defined in embedded space. For the first step, we use an embedding based on time-respecting random walks over the temporal graph. Such walks are known to depend on and to encode important structural and temporal properties of time-varying graphs [70, 9, 58, 43, 59]. For the second step, we proposed two possibilities for the distance definition: if a mapping is known between the nodes of the graphs to be compared, we consider a distance definition that leverages such mapping; in the more general case, when such a mapping is unavailable, we put forward a definition that makes it possible to compare graphs with arbitrarily different sizes (numbers of nodes). In both cases, since the size of the embedding matrix we use does not depend on the graph’s temporal span, it is possible to embed temporal graphs with different durations in the same embedding space and thus compute a distance between them. The evaluation of our approach on both synthetic and empirical data shows that the proposed distance is sensitive to structural differences (e.g., degree distribution, clustering coefficient, or presence of communities) as well as to temporal differences (e.g., burstiness or node activity patterns) of the temporal graphs being compared.
@@
The methodology we developed is very general and customizable, as several alternative choices are possible for both steps of our approach, namely the chosen embedding method and distance. The specific choices we study here are motivated by two considerations: i) a theoretical one, given the importance of time-respecting random walks in encoding information on temporal graphs, and thus in generating distributed representations of temporal graphs [70, 9, 58, 59]; ii) a numerical one, which led us to choose the EDRep algorithm of Ref. [60] because of its efficiency and scalability. Our evaluation results support these choices a posteriori.
@@
Some limitations of our approach are worth mentioning. First, the part of our evaluation based on empirical data focused on a limited set of temporal graphs, namely temporal networks describing the close-range proximity of humans. However, these temporal graphs are known to feature a broad variety of representative structural and temporal features, such as time-varying community structures, burstiness of edge activity, fat-tailed distributions of interaction durations, and more  [13, 15]. Second, we only considered the cases of fully matched graphs (a bijective relation between nodes) or no known matching. An interesting intermediate situation would be partially matched graphs, in which only a subset of nodes are matched across the two graphs. Tackling this challenging case would yield an interesting extension of our work. Finally, our unmatched distance allows us to compare graphs with sizes (number of nodes) that potentially differ even by orders of magnitude. Even if our definition can deal with such extreme cases, we believe that this kind of comparison calls for a more profound question on what it means to compare entities that differ so much. This case might require tailored definitions of distance that leverage domain-specific knowledge.
@@
Defining a distance between temporal graphs opens the door to a wealth of potential applications, of which our evaluation only offers possible examples. On the one hand, such a distance can support data analysis of temporal graph data by extending the procedure of Figure 3 to a richer set of randomizations [50]. Indeed, as discussed, the inability to distinguish pairs of randomizations of a given temporal graph can be regarded as an invariance property of the graph itself. On the other hand, computing a distance between graphs can provide a crucial tool for the validation of generative models: the validation of generative models is often limited to comparing a set of statistical properties [41], while our distance definition enables us to carry out a comparison at the global level. In addition, our distance could underpin the very process of generating synthetic temporal graphs by enabling approaches such as GANS [71] that depend on a global distance and have proved extremely powerful in generating realistic data [72, 73]. In particular, generating synthetic human proximity data with realistic topological and temporal properties could be used to better simulate infectious disease spread and other phenomena of interest for public health research [74, 75, 76, 77, 78, 79]. Finally, we know that empirical (temporal) graphs are hard to anonymize [80, 81, 78]. To minimize the risk of node re-identification, performing some perturbation operations on the graph might be necessary before making it publicly available. However, such operations risk destroying essential patterns and information of empirical data. The distance we introduced could help tackle the trade-offs between re-identification risk and information loss by better quantifying the latter.